OpenStudy (anonymous):

limits- based on your understanding of of the idea of limit, explain why g(x)=(sqroot(100+x)-sqroot(100-x))/x = G(x)=(sqroot(100+x^n)- sqroot(100-x^n))/x^n

OpenStudy (anonymous):

\[\sqrt{100+x}-\sqrt{100-x})/x = \sqrt{100+x ^{n}}-\sqrt{100-x ^{n}}\]

OpenStudy (anonymous):

that last equation should be be divided by x^n as well